### v1

# Resource Allocation for Twin Maintenance and Computing Task Processing in Digital Twin Vehicular Edge Computing Network

Abstract
As a promising technology, vehicular edge computing (VEC) can provide computing and caching services by deploying VEC servers near vehicles. However, VEC networks still face challenges such as high vehicle mobility. Digital twin (DT), an emerging technology, can predict, estimate, and analyze real-time states by digitally modeling objects in the physical world. By integrating DT with VEC, a virtual vehicle DT can be created in the VEC server to monitor the real-time operating status of vehicles. However, maintaining the vehicle DT model requires ongoing attention from the VEC server, which also needs to offer computing services for the vehicles. Therefore, effective allocation and scheduling of VEC server resources are crucial. This study focuses on a general VEC network with a single VEC service and multiple vehicles, examining the two types of delays caused by twin maintenance and computational processing within the network. By transforming the problem using satisfaction functions, we propose an optimization problem aimed at maximizing each vehicle’s resource utility to determine the optimal resource allocation strategy. Given the non-convex nature of the issue, we employ multi-agent Markov decision processes to reformulate the problem. Subsequently, we propose the twin maintenance and computing task processing resource collaborative scheduling (MADRL-CSTC) algorithm, which leverages multi-agent deep reinforcement learning. Through experimental comparisons with alternative algorithms, it demonstrates that our proposed approach is effective in terms of resource allocation.
Index Terms:
Digital twin, Twin maintenance,Vehicular edge computing, Resource allocation
I Introduction
With the rapid advancement of fifth generation (5G) technology, there is a growing presence of vehicular applications and multimedia services in areas such as autonomous driving, navigation, high-definition video, and more. These advancements aim to enhance the overall driving experience [1]-[12] but also lead to an increase in the number of vehicular computing tasks that need to be addressed. However, the storage capacity and computing capabilities of vehicles are often insufficient to handle such demands, posing a challenge for real-time processing of computationally intensive tasks [13]-[20]. In response to this challenge, vehicle edge computing has emerged as a promising solution. By deploying a VEC server at roadside locations, it becomes possible to offload computational tasks from vehicles to these servers for processing and then return the results efficiently [21]-[32].
Although VEC can provide computing services for vehicles, it also faces issues such as vehicle mobility and environmental dynamics during implementation[33]. Digital twin (DT), as a promising and emerging innovation, facilitates the creation of virtual models that accurately represent physical objects [34]-[36]. With advancements in 5G, edge computing, artificial intelligence, and related technologies, the capabilities of DT are continually enhancing. It now goes beyond one-way mirror simulation to enable two-way information interaction [37]. This bidirectional mapping offers a wealth of information for VEC network, enabling feature extraction and prediction of physical vehicles and their surrounding environment [38],[39]. Specifically, vehicles can transmit their data to the server using Vehicle-to-Everything (V2X) technology, allowing them to establish digital replicas on the server based on historical data, leading to the creation of DT models [40].
The combination of DT and VEC can not only collect real-time operational data of vehicles in the VEC network, but also carry out the real time control and change of the status of vehicles[41]. Real-time vehicle information can be accessed and compared with historical data to identify and address potential problems or risks in vehicle operation proactively[42]. However, this integration also presents challenges, particularly in terms of resource management. Establishing and maintaining DT models requires computing resources for information synchronization, while processing vehicle computing tasks on the server further consumes resources. Therefore, it is crucial to explore how to develop reasonable resource allocation strategies to prevent delays in task execution.
In this paper, we consider a DT scenario for mobile edge network with multiple vehicles associated with a single VEC server and propose a multi-agent based twin maintenance and computing task processing resource collaborative scheduling algorithm (MADRL-CSTC) [^1] The source code has been released at: https://github.com/qiongwu86/Resource-allocation-for-twin-maintenance-and-computing-tasks-in-digital-twin-mobile-edge-network., which maximizes the resource utility of computing resource allocation while ensuring the time limit requirements of twin maintenance and vehicle computing tasks. The main contributions of this article are summarized as follows:
  • We consider a general scenario of a DT mobile edge network with multiple vehicles and a single VEC server. We first analyzed the two types of delays caused by simultaneous vehicle digital twin maintenance and computational task processing in this scenario. Then, describe these two types of delays as the resource utility under the current server computing resource allocation strategy using a satisfaction function. Afterwards, considering various constraints, the optimization problem of maximizing the resource utility of each vehicle is proposed.
  • Due to the non convexity of the optimization problem, taking into account the total server resources in the scenario, the deadline for twin maintenance and computing tasks, the maximum transmission power of vehicles, interference between vehicles and servers, and vehicle mobility, we reconstructed the problem using a multi-agent Markov decision process (MDP). Specifically, each vehicle is modeled as an intelligent agent interacting with the environment, i.e. other intelligent agents. Considering the problem itself, we set the reward to the resource utility of each vehicle.
  • In order to explore the optimal allocation strategy for maximizing the utility of vehicle resources, we propose the MADRL-CSTC algorithm. Specifically, we utilized centralized training, distributed execution, and actor-critic (AC) framework, while designing estimation and target AC neural networks for policy iteration and value iteration, respectively. We also use $\epsilon$-greedy algorithm to balance the relationship between exploration and utilization. By comparing it with other algorithms through extensive experiments, the superiority and effectiveness of our algorithm can be verified.
The remaining part of this article is organized as follows. Section II reviews the related work. Section III introduces the system model and analyzes two types of delays caused by twin maintenance and computational task processing, and then uses satisfaction functions to formulate optimization problems. In section IV, we reconstruct the problem using multi-agent MDP and propose the MADRL-CSTC algorithm. Section V presents simulation results to demonstrate the effectiveness of our proposed algorithm. Section VI draws conclusions.
II Related work
In this section, we first review the relevant work on DT, and then investigate the existing work on the combination of VEC and DT.
II-A Digital Twin
In recent years, there has been a significant amount of research on DTs. In [43], Meysam et al. first modeled the BS energy-saving problem as a Markov Decision Process (MDP) to address the significant delay caused to users by not entering sleep mode at the correct time during energy-saving. They also encapsulated the dynamics of the studied system using a DT model and used DT to estimate the risk of decision-making of BS entering sleep mode in advance. In [44], Liao et al. addressed the lack of effective coordination in the river area of vehicle ramps and established a DT model for drivers and vehicles, synchronizing data with real-world vehicles. The data was processed using the DT model and the results were returned to both vehicles and drivers, thus solving the sustainability problem of ramp merging. In [45], He et al. integrated DT and mobile edge computing into the federated learning framework of heterogeneous cellular networks, and assisted local training of user equipment to reduce latency by deploying DT network models at macro base stations. In [46], Liu et al. combined DT with edge collaboration and proposed DT assisted mobile users for task offloading and modeled it as MDP. Then, they decomposed it into two sub models and solved them using decision tree algorithm and dual deep qlearning. In [47], Lu et al. integrated DTs with edge networks, proposed DT edge networks, and used deep reinforcement learning to solve the placement and transfer problems of DTs caused by network dynamics and topology changes.
Figure 1: Digital twin mobile edge network.
II-B Digital Twin Vehicular Edge Computing
Recently, there have been some studies combining DTs with VEC. In [33], Dai et al. proposed an adaptive DT VEC network that utilizes DT caching content and designs an offloading scheme based on the DRL framework to minimize offloading latency. In [48], Zhang et al. combined DTs with artificial intelligence into the VEC network, guiding vehicles to aggregate edge services by deploying digital twins in road side units (RSUs) to minimize offloading costs. In [49], Liao et al. considered that in mixed traffic scenarios, the planned operation of autonomous vehicles (CAVs) may be affected by human driven vehicles (HDVs), and developed a driver DT model that was deployed on the server to help CAVs predict the possible lane changing behavior of surrounding HDVs. In [50], Zhang et al. proposed a socially aware vehicle edge caching mechanism, which constructs a vehicle social relationship model by deploying a digital twin model of edge caching systems on RSUs and designs an optimal caching scheme based on DRL. In [51], Zheng et al. proposed that vehicles in heterogeneous vehicle networks need to improve communication efficiency by selecting different networks, while also facing situations of uneven vehicle distribution and dynamic networks. They established a prediction model in DT to predict the waiting time for vehicles to connect to the network and return the results to the vehicles for decision-making.In [52], Zhao et al. established a vehicle DT model in RSUs to learn the global information of the vehicle, thereby assisting clustering algorithms in reducing the scope of vehicle task offloading and achieving the goal of offloading prediction. In [53], Sun et al. proposed the Digital Twin Edge Network (DITEN) and established DT models for edge servers and the entire MEC system to provide auxiliary services for mobile user business offloading, thereby minimizing offloading latency. In [54], Li et al. integrated unmanned aerial vehicles (UAVs) into VEC and established DT models for RSUs, UAVs, and vehicles in the central controller to manage UAV and RSU resources and assist in offloading vehicular tasks.Although establishing a DT model of vehicles in the servers of the VEC network can help vehicles perform better task offloading, edge business aggregation, or execute specific traffic behaviors, at the same time, the maintenance of the vehicle DT model will also consume server resources, which also needs to be considered.
As mentioned above, there has been no research on resource allocation issues in the coexistence of vehicle twin maintenance and VEC server services for vehicles.
III System model
As shown in Figure 1, we consider a DT scenario of Mobile Edge Network (DTMEN) with $N$ vehicles and a server equipped base station. For clarity, let $N\stackrel{{\scriptstyle\Delta}}{{=}}\{1,2,3,...,N\}$ represent the number of vehicles driving on the lane, and its initial position follows the Poisson clustering process. Specifically, these vehicles will travel in one direction along the lane and generate computing tasks during their journey. Due to their limited computing resources, the vehicles will connect to the base station (BS) through the V2I link using a cellular interface (Uu) to offload the computing tasks to the server for processing. The completed results will be returned to the vehicles. In addition, DT models are established for these vehicles on the server to obtain real-time operational status of the vehicles. Meanwhile, in order to maintain the digital twin models of these vehicles, the server needs to synchronize information with the vehicles to ensure that the models can accurately reflect the driving conditions of the vehicles.
III-A Vehicle Movement Model and Channel Model
III-A1 Vehicle Movement
Similar to [55], a spatial orthogonal coordinate system is established with the base station as the origin in figure 2. The positive direction of the x-axis is the direction of the vehicle traveling along the lane, that is, east, the positive direction of the y-axis is south, and the positive direction of the z-axis is along the direction of the base station antenna. Therefore, the coordinate of vehicle $i(i\in\{1,2,3...,N\})$ located on lane j at time slot t can be represented as $PO_{i}(t)$
$PO_{i}(t)=(X_{i,j}(t),Y_{i,j}(t),0),i\in\{1,2,3...,N\},$ (1)
where $X_{i,j}(t)$ is the horizontal ordinate of vehicle $i$ driving on lane $j$ in time slot $t$, $Y_{i,j}(t)$ is the ordinate of vehicle $i$.
When the value of each time slot $\tau$ is small enough, the position of the vehicle in each time slot can be approximately considered constant[56]. At the same time, the position of the vehicle in the current time slot $t$ is related to the position in the previous time slot $t-1$. Therefore, horizontal ordinate $X_{i,j}(t)$ of vehicle $i$ driving on lane $j$ in time slot $t$ can be further represented as
$X_{i,j}(t)=X_{i,j}(t-1)+\tau v_{i},$ (2)
where $\tau$ is the duration of each time slot and $v_{i}$ is the driving speed of vehicle $i$.
Due to vehicle $i$ drives on lane $j$, we let $w_{0}$ represents the width of each lane, and $L_{0}$ represents the distance between the first lane and BS.Therefore, the ordinate $Y_{i,j}(t)$ of vehicle $i$ can be written as
$Y_{i,j}(t)=L_{0}+jw_{0},j\in(0,1,2,3...J-1),$ (3)
Figure 2: Vehicle movement model.
III-A2 Channel Model
We use Orthogonal Frequency Division Multiple Access (OFDMA) in DTMEN for communication between vehicles and base stations, so we don’t consider interference between vehicles[57][58]. At the current time slot $t$, the channel gain between vehicle $i$ and BS, $g_{i,B}^{t}$ is modeled as [59]
$g_{i,B}^{t}=|h_{i,B}^{t}|^{2}v_{i,B}^{t},t=0,1,2,...,$ (4)
where $v_{i,B}^{t}$ is the large-scale fading component between vehicle $i$ and BS, which is composed of path loss and shadow distribution. $h_{i,B}^{t}$ is the small-scale path fading component between vehicle $i$ and BS at time slot $t$, which follows a circularly symmetric complex Gaussian distribution with unit variance. We use a first-order Gaussian Markov process to describe small-scale path fading. Therefore, the update of $h_{i,B}^{t}$ can be represented as
$h_{i,B}^{t}=\kappa h_{i,B}^{t-1}+\frac{1}{(\sqrt{(PO_{i}(t)-P_{B}(t))^{2}})^{\beta}},$ (5)
where $\kappa$ is correlation coefficient, $P_{B}(t))^{2}$ is the coordinates of the base station, $\beta$ is a path loss index.
III-B Communication Model and Computing Model
III-B1 Communication Model
As shown in Figure 1, the communication between the vehicle and the base station includes information updates for vehicle twin maintenance, uplink communication for task transmission with the same server, and downlink communication for returning calculation results to the vehicle. It should be noted that since the size of the results processed by the digital twin edge server is much smaller than before, and the communication rate of the downlink is much larger, Therefore, we ignore the delay caused by downlink communication. According to Shannon’s theorem, we can calculate the uplink communication rate between the vehicle and the base station as follows[60]
$r_{i,B}^{t}=W\log(1+\frac{p_{i}g_{i,B}^{t}}{\sigma^{2}}),$ (6)
where $W$ is the bandwidth for wireless transmission between vehicle $i$ and BS, $p_{i}$ is the transmit power of vehicle $i$, $\sigma^{2}$ is the noise between vehicle $i$ and BS.
III-B2 Computing Model
After establishing the digital twin of vehicle $i$, in order for the digital twin model to accurately reflect the current situation of the vehicle, the digital twin of vehicle $i$ in the VEC server needs to be synchronized with the information of vehicle $i$ in the physical space[38]. We let $\Gamma_{i}(t)=\{{D_{i}^{dt}(t),C_{i}^{dt}(t),T_{i}^{dt}}\}$ represents the information of vehicle $i$ that needs to be synchronously updated at the current time slot $t$, where $D_{i}^{dt}(t)$ is the size of the information of vehicle $i$ that needs to be updated at the current time slot $t$, $C_{i}^{dt}(t)$ represents the CPU frequency required to update the unit size of vehicle $i$ information, and $T_{i}^{dt}$ is the maximum delay limit for updating vehicle $i$ information. The updated information needs to be transmitted from vehicle $i$ to the server equipped with the base station first, and then the server will allocate computing resources to maintain the digital twin of vehicle $i$. Here, according to the hosting restrictions of [61], the maintenance of digital twins on vehicle $i$ is also related to the total number of digital twins it needs to maintain on its server. In addition, we record the total amount of computing resources on the server as $F$, and the computing resources requested by vehicle $i$ for twin maintenance as $f_{i}^{dt}$. Therefore, the time $t_{i}^{dt}$ required to complete the digital twin maintenance of vehicle $i$ can be calculated as
$t_{i}^{dt}=\frac{D_{i}^{dt}(t)}{r_{i,B}^{t}}+\frac{D_{i}^{dt}(t)C_{i}^{dt}(t)}{f_{i}^{dt}}N,$ (7)
Vehicles may generate computing tasks while driving, and due to their insufficient computing power, they may choose to offload the computing tasks to edge servers for processing. We represent the computing task generated by vehicle $i$ at time slot $t$ as $\zeta_{i}(t)=\{{D_{i}^{tk}(t),C_{i}^{tk}(t),T_{i}^{tk}}\}$, where $D_{i}^{tk}(t)$ is the computing task size generated by vehicle $i$ at time slot $t$, $C_{i}^{tk}(t)$ is the CPU frequency required to compute unit size tasks, and $T_{i}^{tk}$ is the maximum delay limit for processing computing tasks. Similar to the previous maintenance of digital twins, the vehicle $i$ will first offload the computing task to the server, and then the server will allocate computing resources to handle the computing task. Therefore, the time $t_{i}^{tk}$ required to complete the calculation task, which can be calculated as
$t_{i}^{tk}=\frac{D_{i}^{tk}(t)}{r_{i,B}^{t}}+\frac{D_{i}^{tk}(t)C_{i}^{tk}(t)}{f_{i}^{tk}},$ (8)
where $f_{i}^{tk}$ represents the computing resources requested by vehicle $i$ for processing computing tasks transmitted to the server.
It should be noted that the maintenance and computing tasks of the digital twin of vehicle $i$ are carried out simultaneously, so the sum of the computing resources $f_{i}^{dt}$ allocated for the maintenance of the digital twin of vehicle $i$ and the computing resources $f_{i}^{tk}$ allocated for processing the computing tasks generated by vehicle $i$ can’t exceed the total computing resources $F$ of the server, that is, $\Sigma_{1}^{N}f_{i}^{dt}+f_{i}^{tk}\leq F$.
III-C Optimization Problem
Due to the fact that vehicle $i$ simultaneously performs digital twin maintenance and offloading of computing tasks at time slot $t$, VEC server will allocate computing resources for both, namely $f_{i}^{dt}$ and $f_{i}^{tk}$, respectively. By using different computing resource allocation strategies, different information synchronization delays and task computation delays can be obtained, which in turn affects the maintenance delay $t_{i,B}^{dt}$ of digital twins and the processing delay $t_{i,B}^{tk}$ of computing tasks. According to [62], we define a satisfaction function $Q$ to demonstrate the satisfaction level of digital twin maintenance delay and computational task processing delay for vehicle $i$ under a certain resource allocation strategy $\omega_{i}$, respectively. Therefore, the satisfaction function $Q_{i}^{dt}(\omega_{i})$ for the maintenance delay of vehicle $i$ digital twin and the satisfaction function $Q_{i}^{tk}(\omega_{i})$ for the processing delay of the computational task for vehicle $i$ can be respectively represented as
$Q_{i}^{dt}(\omega_{i})=1-\ln(1+\frac{t_{i}^{dt}}{T_{i}^{dt}}),$ (9)
$Q_{i}^{tk}(\omega_{i})=1-\ln(1+\frac{t_{i}^{tk}}{T_{i}^{tk}}),$ (10)
Through (9) and (10), we can evaluate the resource utility $U_{i}(\omega_{i})$ obtained by vehicle $i$ under a resource allocation strategy $\omega_{i}$
$U_{i}(\omega_{i})=\rho Q_{i}^{dt}(\omega_{i})+(1-\rho)Q_{i}^{tk}(\omega_{i}),$ (11)
where $\rho$ is the weighting factor with $10<\rho<$ 1 between twin maintenance and computational task processing, used to measure the importance of the two tasks.
Our goal is to maximize the resource utility $U$ obtained by each vehicle while maintaining a fixed total server resource and meeting the maximum latency conditions for twin maintenance and computing tasks. Therefore, the optimization problem under consideration can be formulated as
$P1:{\rm{}}\mathop{\max}_{i\in N}\quad U_{i}(\omega_{i})\$ (12a)
${\rm{}}s.t.\quad{\rm{}}t_{i}^{dt}\leq T_{i}^{dt},\$ (12b)
$\qquad{\rm{}}t_{i}^{tk}\leq T_{i}^{tk},\$ (12c)
$\qquad\qquad\quad\sum_{i=1}^{N}f_{i}^{dt}+f_{i}^{tk}\leq F,\$ (12d)
$\qquad\qquad\quad p_{i}\leq p_{i}^{max},\forall i\in N,\$ (12e)
where $p_{i}^{max}$ is the maximum transmit power of vehicle i. Constraint (12b) requires that the maintenance time of digital twins cannot exceed the maximum delay tolerance. Constraint (12c) requires that the processing time of computing tasks cannot exceed the maximum delay tolerance. Constraint (12d) requires that the sum of computing resources consumed by all vehicles cannot exceed the server’s own computing resource capacity. Constraint (12e) means that the transmission power of each vehicle cannot exceed the maximum value.
IV Proposed solution
Each vehicle in the scene will simultaneously perform digital twin maintenance and offloading of computing tasks to the server. The computing resources requested by a vehicle to the server to complete these two tasks will also affect the resources obtained by other vehicles, thereby affecting the final utility. In addition, the computing resources requested by each vehicle can be either integer or real variables. Obviously, our optimization problem is a mixed integer nonlinear programming problem, which is NP-hard and cannot be completed in polynomial time using traditional methods. Therefore, we transformed the problem into a multi-agent MDP and proposed a MARL-CSTC algorithm to solve it, while analyzing the complexity of the algorithm.
IV-A Multi-agent MDP transformation of problems
The multi-agent MDP problem can be described by using a five tuple $(\mathcal{N,S,A,P,R})$, where $\mathcal{N,S,A,P,R}$ in the tuple represent the set of agents, state space, action space, state transition probability, and reward function, respectively.
IV-A1 Agent Set $\mathcal{N}$
In order to achieve the goal of maximizing vehicle utility, each vehicle will act as an agent to learn resource allocation for digital twin maintenance and computing tasks. So, we let the agent set $\mathcal{N}=\{1,2,3,...N\}$
IV-A2 State Space $\mathcal{S}$
The state $s_{n}(t)$ of agent $n$ at moment $t$ can be described by updating information, computing tasks, vehicle position, speed, and channel gain, which can be represented as $s_{n}(t)=\{\Gamma_{n}(t),\zeta_{n}(t),PO_{n}(t),v_{n}\}$. The state of the entire system $s(t)$ can be expressed as $s(t)=\{s_{n}(t)\}_{N}$.
IV-A3 Action Space $\mathcal{A}$
The behavior $a_{n}(t)$ taken by the agent $n$ includes computing resources for the maintenance of the digital twin of agent $n$, $f_{n}^{dt}$, as well as resources for the processing of computing tasks of agent $n$, $f_{n}^{tk}$. Therefore, the behavior $a_{n}(t)$ taken by agent $n$ can be expressed as $a_{n}(t)=\{f_{n}^{dt},f_{n}^{tk}\}$. The action space $a(t)$ for all agents can be represented as $a(t)=\{a_{n}(t)\}_{N}$.
IV-A4 State Transition Probability $\mathcal{P}$
The state transition probability $p_{n}(t)$ describes the probability that at each decision epoch $t$, the agent $n$ executes action $a_{n}(t)$ in the current state $s_{n}(t)$ and then transfers to the next state $s_{n}(t+1)$, that is, $p_{n}(s_{n}(t+1);s_{n}(t),a_{n}(t))$.
IV-A5 Reward Function $\mathcal{R}$
The reward function will return the corresponding reward value or a penalty value when the agent $n$ takes action in a given state. By considering the optimization problem itself and its constraints, we let the Utility $U$ serve as the reward function, that is, the reward function $r_{n}(t)$ of agent $n$ is
$r_{n}(t)=\rho Q_{n}^{dt}(\omega_{n})+(1-\rho)Q_{n}^{tk}(\omega_{n}),$ (13)
on this basis, the long-term discount reward we can receive is
$R_{n}(t)=\sum_{t_{0}}^{t}{\gamma_{n}r_{n}(t)},$ (14)
where $t_{0}$ is the previous epoch, $\gamma_{n}$ is the discount factor, with a range of values $\gamma_{n}\in[0,1]$, which represents the degree to which past rewards have an impact on the rewards at the current epoch $t$.
Figure 3: Structure of MARL-CSTC
IV-B Algorithm Design
As shown in Figure 3, for the structure of MARL-CSTC algorithm, actor-critic framework is employed, where the actor is used to generate the actions that the agent needs to perform, while the critic is responsible for guiding the actor network to generate better actions. Here, the actor consists of two parts, where the estimation actor network is responsible for training, and the target actor network is responsible for executing the action. Similarly, critic also includes two parts, namely the estimation critic network and the target critic network, both of which are used to evaluate the actions of agents. For this reason, the actor network adopts a policy based deep neural network, while the critic uses a value based deep neural network. In addition, considering the dynamic changes in the environment, we adopt a centralized training and distributed execution strategy, where the critic is centrally trained by the server and the actor is executed by each vehicle in a distributed manner.
IV-B1 actor
We construct the actor network into three layers: input layer, fully connected layer, and output layer. The fully connected layer is composed of three hidden layers and one softmax layer. The first two hidden layers use the rectified linear unit (ReLu) function as the activation function, and the third hidden layer uses the tangent (tanh) function as the activation function. By constructing an actor network in this way, the input state can be transformed into all possible actions, namely resource allocation strategies. Since the actor network is divided into two parts, we will explain the estimation actor network and the target actor network separately.
For the estimation actor network of agent $n$, its input is the current state $s_{n}(t)$, including the synchronization information required for twin maintenance $\Gamma_{n}(t)$, computational tasks generated by vehicles $\theta_{n}(t)$, vehicle position $PO_{n}(t)$, and vehicle speed $v_{n}$. The current state is input into the fully connected layer, processed by three hidden layers, and output as the probability of all possible actions. After passing through the softmax layer, the total probability of all actions is set to 1. Finally, agent $n$ selects one of them as the final resource allocation strategy to be executed.
Similar to estimation actor network, the input of target actor network is the next state $s_{n}(t+1)$, and the corresponding output is the next action $a_{n}(t+1)$ processed by the fully connected layer. It is should be noted that although the structure of target actor network and estimation actor network are with the similar structure, their parameters $\theta^{{}^{\prime}}(\pi_{n})$ and $\theta(\pi_{n})$ are different respectively, where the former is the target actor network parameter and the latter is the estimation actor network parameter.
IV-B2 critic
Similar to the actor network, the structure of the critic network includes an input layer, a fully connected layer, and an output layer. The difference is that only the first three layers in the fully connected layer have activation functions, namely Relu layer, Relu layer, and Tanh layer.
For the estimation critic network of agent $n$, its input is the state and actions of all agents at the current epoch, namely $\mathcal{S}$ and $\mathcal{A}$. After the current input is processed by the fully connected layer, the Q value is obtained. For each decision epoch $t$, the Q value of agent $n$ can be defined as $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})=\mathbb{E}_{\pi}[R_{n}(t);s(t),a(t);\theta_{Q_{n}}]$.
For the target critic network of agent $n$, similar to the estimation critic network, the input is the state and action of all agents at the next epoch, and after being processed by the fully connected layer, the corresponding $Q$ value $Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta_{Q_{n}}^{{}^{\prime}})$ is obtained. It should be noted that the structure of the target critic network is the similar to that of the estimation critic network, but their parameters are different, i.e. $\theta^{{}^{\prime}}_{Q_{n}}$ and $\theta_{Q_{n}}$.
Algorithm 1 Training Stage of MARL-CSTC Algorithm.
0:  $N,D_{n}^{dt},C_{n}^{dt},T_{n}^{dt},D_{n}^{tk},C_{n}^{tk},T_{n}^{tk},v_{n},g_{n,B}^{t}$ for $n=1,...,N$;
0:  $\theta_{\pi_{n}}^{{}^{\prime}}$ for $n=1,...,N$
1:  Initialize discount factor $\gamma$ and parameter update rate $\eta$;
2:  Random initialize the $\theta_{Q_{n}},\theta_{Q_{n}}^{{}^{\prime}},\theta_{\pi_{n}},\theta_{\pi_{n}}^{{}^{\prime}}$ for $n=1,...,N$;
3:  for $k$ eposide from 1 to $K$ do
4:     for $n$ agent from 1 to $N$ do
5:        Initialize $s_{n}(t)$;
6:        Input $s_{n}(t)$ to estimation actor network, and get $a_{n}(t)=\pi_{n}(s_{n}(t);\theta_{\pi_{n}})$;
7:        Execute $a_{n}(t)$ based on $s_{n}(t)$, obtain $r_{n}(t)$ and transfer to $s_{n}(t+1)$;
8:        Store $(s_{n}(t),a_{n}(t),r_{n}(t),s_{n}(t+1))$ as an experience in the Replay Buffer;
9:        Input $\mathcal{S}$ and $\mathcal{A}$ to estimation critic network and compute $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$;
10:        Input $\mathcal{S^{{}^{\prime}}}$ ,$\mathcal{A^{{}^{\prime}}}$ to target critic network and compute $Q_{n}^{{}^{\prime}}(\mathcal{S^{{}^{\prime}}},\mathcal{A^{{}^{\prime}}};\theta_{Q_{n}}^{{}^{\prime}})$;
11:        Calculate Q value, temporal difference $\delta$ and loss function $L(\theta_{Q_{n}})$;
12:        Update $\theta_{Q_{n}}$ by stochastic gradient descent;
13:        Input $s_{n}(t)$ to estimation actor network and obtain $a_{n}(t)=\pi_{n}(s_{n}(t);\theta_{\pi_{n}})$;
14:        Input $s_{n}(t+1)$ to target actor network and obtain $a_{n}(t+1)=\pi_{n}^{{}^{\prime}}(s_{n}(t+1);\theta_{\pi_{n}}^{{}^{\prime}})$;
15:        Update $\theta_{\pi_{n}}$ by gradient descent;
16:        Update $\theta_{Q_{n}}^{{}^{\prime}}$ and $\theta_{\pi_{n}}^{{}^{\prime}}$ , respectively;
17:     end for
18:  end for
IV-C Algorithm Training
The centralized training of the MARL-CSTC algorithm is implemented on the server, as summarized in Algorithm 1. Specifically, since the server manages the critic network of each agent, the status and actions of all agents can be obtained, meaning that the information of all agents is observable. In the centralized training phase, the server first obtains the states and actions of all agents. By utilizing this information, the server can train the estimation critical network for each agent, thereby achieving the goal of maximizing the Q value. For a single agent $n$, its Q-value $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$ can be updated through the Bellman equation[63], which is given by $Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})=R_{n}(t)+\gamma_{n}\max Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})$. The time difference error can be calculated as $\delta=Q_{n}^{{}^{\prime}}(\mathcal{S}^{{}^{\prime}},\mathcal{A}^{{}^{\prime}};\theta^{{}^{\prime}}_{Q_{n}})-Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})$. Therefore, the loss function can be expressed as $L(\theta_{Q_{n}})=\mathbb{E}(\delta^{2})$. To minimize the loss function $L(\theta_{Q_{n}})$, we update the parameters $\theta_{Q_{n}}$ by using a stochastic gradient descent algorithm, which is represented as $\nabla_{\theta_{Q_{n}}}L(\theta_{Q_{n}})=\mathbb{E}(2\delta\nabla_{\theta_{Q_{n}}}Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}}))$.
Unlike critic networks, due to the distributed execution of actor networks and the need for agents to take actions based on local observations, actor networks are deployed on each vehicle. The parameters of the actor network are updated through gradient descent
$\nabla_{\theta_{\pi_{n}}}L(\theta_{\pi_{n}})\approx\mathbb{E}[\nabla_{\theta_{\pi_{n}}}\log\pi_{n}(s_{n}(t);\theta_{\pi_{n}})Q_{n}(\mathcal{S},\mathcal{A};\theta_{Q_{n}})],$ (15)
where $\pi_{n}(s_{n}(t);\theta_{\pi_{n}})$ represents the strategy for executing actions in the current state.
To ensure the stability of the entire training process, we adopt a soft update approach to update the parameters of the target actor network and target critic network. The specific update methods are $\theta_{Q_{n}}^{{}^{\prime}}=\eta\theta_{Q_{n}}+(1-\eta)\theta_{Q_{n}}^{{}^{\prime}}$ and $\theta_{\pi_{n}}^{{}^{\prime}}=\eta\theta_{\pi_{n}}+(1-\eta)\theta_{\pi_{n}}^{{}^{\prime}}$, respectively, where $\eta$ is parameter update rate, while $\eta\in[0,1]$.
Algorithm 2 Executing Stage of MARL-CSTC Algorithm.
0:  $N,\theta_{\pi_{1}}^{{}^{\prime}}$, $\theta_{\pi_{2}}^{{}^{\prime}}$,…, $\theta_{\pi_{N}}^{{}^{\prime}}$;
0:  $a_{n}(t)$ for $n=1,...,N$
1:  for agent $n$ from 1 to $N$ do
2:     Input $\theta_{\pi_{n}}^{{}^{\prime}}$ to the target actor network of agent $n$;
3:     Agent $n$ observe the environment to get current state $s_{n}(t)$;
4:     Choose a probability $Pr_{n}\in[0,1]$;
5:     if $Pr_{n}\leq\epsilon$ then
6:        Randomly select an action $a_{n}(t)\in\mathcal{A}$;
7:     else
8:        Compute the value of action by $a_{n}(t)=\pi_{n}^{{}^{\prime}}(s_{n}(t);\theta_{\pi_{n}}^{{}^{\prime}})$;
9:        Select the action with the highest Q value;
10:     end if
11:  end for
IV-D Algorithm Execution
After the centralized training is completed, distributed actions need to be executed on each vehicle based on local observations. Specifically, agent $n$ downloads the results of centralized training and inputs them into the actor network. Then, agent $n$ observes and obtains the local state $s_{n}(t)$, and executes action $a_{n}(t)$ based on policy $\pi_{n}$ to obtain a reward $r_{n}(t)$. In addition, due to a lack of experience in the early stages of execution, agent $n$ randomly selects actions to explore. When experience is sufficient, it will execute actions to maximize rewards. Therefore, in order to achieve a balance between exploration and exploitation, we adopt the $\epsilon$-greedy algorithm, as summarized in Algorithm 2.
IV-E Algorithm Complexity Analysis
Inspired by [37], the computational complexity of algorithms is mainly related to the structure and parameters of neural networks. Due to the fact that both the actor network and the critic network of the MARL-CSTC algorithm use DNN, the calculation of complexity needs to be based on the analysis of DNN. We let $Ln$ represent the number of layers in DNN, and let $G_{l}$ represent the number of neurons in the $l$-th layer. Therefore, we can obtain that the computational complexity of the actor network and the critic network is $\mathcal{O}(X_{a})=\mathcal{O}(X_{c})=\mathcal{O}(\sum_{l=1}^{Ln}G_{l}G_{l+1})$.
Due to our algorithm adopting a centralized training and distributed execution approach, we still need to analyze these two parts separately. For centralized training, each agent has $E$ experiences stored in the experience replayer, and agent $n(n\in{1,2,...,N})$needs to train $K$ iteration cycles. Therefore, the computational complexity of actors and critics is $\mathcal{O}_{a}(X_{a}KE^{N})$ and $\mathcal{O}_{c}(X_{c}KE^{N})$, respectively. For distributed execution, agent n only needs to observe the state and then execute actions through the actor network, so the complexity of the actor is $\mathcal{O}_{a}(X_{a})$.
V Simulation results
In this section, to evaluate the performance of the MARL-CSTC algorithm on the scenario considered in this paper, we conducted simulation experiments and compared it with other algorithms to analyze its superiority.
V-A Experimental Setup
The simulation experiment in this article was implemented using Python 3.11.0 and Pytorch. In terms of environmental settings, the experimental parameters are listed in Table $\mathrm{I}$. The number of lanes $J$ is set to 3, the bandwidth $W$ is set to 150MHZ, the noise powers $\sigma^{2}$ are set to $10^{-11}mW$ and the information required for twin maintenance and the calculation task size generated by the vehicle itself are set to uniformly sample between [1024,1536] bytes. The CPU frequency required to process these two units of Btye is set to $C_{i}^{dt}=C_{i}^{tk}=0.25$MHZ/Byte, with a maximum delay tolerance of 0.5s. We set the channel correlation coefficient $\kappa$ to 0.2, the path loss index $\beta$ to 3, and set the weight $\rho$ to 0.5, which indicates that twin maintenance and vehicular computing task processing are equally important. In terms of DNN, for actor networks, the number of neurons in the first two hidden layers of the three fully connected layers is 300 and 100, respectively, and the number of neurons in the third hidden layer is the action dimension. For the critic network, the number of neurons in the three hidden layers is 300, 100 and 1, respectively. During the training phase, in terms of learning rate, the actor network has a learning rate of $lr_{a}=2\times 10^{-4}$, the critic network has a learning rate of $lr_{c}=10^{-3}$, and a discount factor $\gamma_{n}=0.95$.
TABLE I: Experimental parameter settings
Parameter Value
$J$ 3
$W$ 150MHZ
$\sigma^{2}$ $10^{-11}mW$
$D_{i}^{dt},D_{i}^{tk}$ [1024,1536]Bytes
$C_{i}^{dt},C_{i}^{tk}$ 0.25MHZ/Byte
$T_{i}$ 0.5s
$v$ [10,15]m/s
$\kappa$ 0.2
$\beta$ 3
$\rho$ 0.5
$\eta$ 0.01
$lr_{a}$ $2\times 10^{-4}$
$lr_{c}$ $10^{-3}$
$\gamma_{n}$ 0.95
Figure 4: Average rewards under different algorithms
V-B Algorithm Comparison
We compared three algorithms, namely Soft Actor-Critic (SAC) algorithm, Proximal Policy Optimization (PPO) algorithm, and Random algorithm.
  • Both SAC algorithm and PPO algorithm can be regarded as single agent algorithms.At each epoch, only one agent updates its actions based on locally obtained information and trained DQN, while the actions of other agents remain unchanged. Single DQN shared among all agents.
  • Random algorithm is the process of first dividing the total computing resources by the number of agents at each decision time t, and then randomly allocating computing resources for twin maintenance and processing computing tasks to the corresponding agents within each average component.
(a) Resource utilization
(b) The ratio of resource conversion to utility
Figure 5: Comparison of resource utilization and the ratio of resource conversion to utility under different algorithms
(a) The impact of the number of vehicles on resource utility:$p_{i}^{max}$=200mW
(b) The impact of maximum transmission power on resource utility:$N$=5
Figure 6: Comparison of resource utility under different algorithms
V-C Performance Evaluation
In Figure 4, we compared the average rewards of MARL-CSTC algorithm with SAC algorithm, PPO algorithm, and random algorithm. As the number of iterations increases, the average rewards of the other three algorithms, except for the random algorithm, will eventually converge to a larger value. Specifically, when the number of iterations is small, due to lack of experience, the total amount of resources requested by the agent from the server may exceed the computational resources of the server itself. In addition, agents may also be punished for violating constraints (16b) and (16c) due to improper resource allocation strategies in the early stage. As the number of iterations gradually increases, the agent is able to explore appropriate resource allocation strategies within the total computational range of the VEC server, resulting in a gradual convergence of reward values. The fluctuation after convergence is due to the dynamic randomness of the environment, which can affect learning. Among the three algorithms, the average converged reward value by the MARL algorithm is greater than that of the SAC algorithm and PPO algorithm. This is because compared to algorithms such as SAC and PPO, our proposed MARL-CSTC algorithm can allocate server computing resources reasonably for twin maintenance and computing task processing based on current number of vehicles, channel resources, transmission power, and other information. The reason why the random algorithm cannot converge is because it first distributes the total resources evenly and then randomly allocates the number of resources, so its average reward value always fluctuates around a certain value, but it is still lower than the average reward value of the final convergence of MARL-CSTC.
Figure 5 shows the resource utilization rate and the ratio of resource conversion to utility under different algorithms and vehicle numbers. From Figure 5 (a), it can be seen that MARL-CSTC achieves the highest resource utilization rate. This is because MARL-CSTC can fully schedule the computing resources, i.e. CPU frequency, of the VEC server based on the twin maintenance and vehicular task requirements of all vehicles. On the contrary, PPO and SAC are based on local information of a single vehicle for resource scheduling and sharing with other vehicles, thus unable to fully utilize server computing resources. Due to the random nature of twin maintenance and vehicular task requirements for individual vehicles, as well as changes in the number of vehicles, the previous resource scheduling plan is often not applicable and needs to be redesigned. Therefore, the resource utilization rates of PPO and SAC will fluctuate. From Figure 5 (b), it can be seen that as the number of vehicles increases, the ratio of computing resources to utility by each algorithm will decrease, while MARL-CSTC can still achieve higher conversion ratios than PPO and SAC. As mentioned earlier, due to its ability to rely on global information, MARL-CSTC is less affected by changes in the number of vehicles. In addition, MARL-CSTC will prioritize meeting the twin maintenance and vehicular task requirements of each vehicle in its specific allocation plan, thus achieving a higher conversion rate.
Figure 6 shows the maximum utility that a single vehicle can achieve under different influencing factors. From Figure 6 (a), it can be seen that when the total number of vehicles is small, a single vehicle can achieve higher resource utility. As the number of vehicles increases, the resource utility of each vehicle will decrease accordingly.This is because an increase in the number of vehicles will require the server to maintain more digital twin models and handle more computing tasks. Therefore, the computing resources allocated to individual vehicles will be reduced. However, the performance achieved under the MARL-CSTC algorithm is still higher than other algorithms. This is because according to Figure 5(a), MARL-CSTC has higher resource utilization than SAC, PPO, and Random, which means that MARL-CSTC has more assignable computing resources. From Figure 6(b), when the total number of vehicles is 5, as the transmission power of the vehicles increases, the utility achieved by each algorithm will also increase. An increase in transmission power means that under the same channel resource conditions, the transmission rate will also increase, which in turn will reduce the possibility of exceeding the time limit for twin maintenance and computing tasks. Therefore, computing resources originally used to meet the time limit conditions can be allocated to achieve higher utility.
(a) Twin maintenance delay
(b) Vehicular computing task processing delay
Figure 7: Comparison of two types of delays under different algorithms:$p_{i}^{max}$=200mW, $N$=5
Figure 7 examines the latency generated by each algorithm while performing twin maintenance and computing task processing simultaneously. We can observe that as the number of vehicles in the environment increases, the latency generated by each algorithm also increases. Specifically, when the number of vehicles is 5, except for the random algorithm, the other three algorithms can achieve lower latency while satisfying time constraints, and the MARL algorithm achieves the minimum latency. As the number of vehicles gradually increases, except for the MARL-CSTC algorithm, which can reduce latency under time constraints in twin maintenance and computational task processing, the other three types of algorithms cannot simultaneously handle these two tasks. According to Figure 5, this is because on the one hand, MARL-CSTC is used for more allocatable computing resources, and on the other hand, it has a higher resource conversion ratio, which prioritizes meeting the time constraints of twin maintenance and onboard tasks, and on this basis, the remaining computing resources are re allocated. In addition, SAC and PPO algorithms tend to favor twin maintenance as the number of vehicles gradually increases, while random algorithms tend to favor computing tasks.
VI Conclusions
This paper considered the server resource allocation problem of twin maintenance and computing task processing simultaneously under edge computing for a digital twin scenario with single server and multiple vehicles, and constructed it as an optimization problem to maximize the resource utility of each vehicle. To address the non convexity of the problem, we first adopted multi-agent MDP to rephrase the problem and proposed the MADRL-CSTC algorithm to obtain the optimal resource allocation. Experimental results have shown that compared with other algorithms, MADRL-CSTC can maximize the resource utility of each vehicle under constraint conditions. The conclusion is summarized as follows:
  • Our method can make real-time decisions on moving vehicles and make decisions on multiple vehicles in the environment simultaneously.
  • Our method provides an effective resource allocation for each vehicle within the total server resource range, under the condition of twin maintenance and computing task deadlines.